I need help on functions... for example like this homework question. f(x)= 3x+5 and g(x)=x^2 my problem is what is the funtion for f(x)+g(x)
Answer by ad_alta(240)

The answer is simple: just add the functions! f(x)+g(x)=[3x+5]+[x^2]=x^2+3x+5